Consider the reaction

Mg(s)+Fe2+(aq)→Mg2+(aq)+Fe(s)

at 57 ∘C , where [Fe2+]= 3.40 M and [Mg2+]= 0.210 M .

Part D Calculate the standard cell potential for Mg(s) Fe2 (aq)Mg2 (aq)Fe(s) Express your answer to three significant figuresPart C What is the value for n? Express your answer as an integer and include the appropriate units (i.e. enter mol for molesReview I Constants Periodic Table Learning Goal: To learn how to use the Nernst equation Consider the reaction The standard rReview I Constants I Periodic Table Part B What is the value for the temperature, T, in kelvins? Express your answer to threehelp
Answer #1

M8 (5) + Fe2+ (aq) → Mg 2+ (aq) + fecs) Here, Mo is oxidised to Mg2+ feat, reduced i Fe. So, the cell be- M9 (5) Mgat 11 featpart 19 570 € = 273+57 - 330K so; T = 330k part At Part @ carnation is written as the Concentrations of products divided by tSo, 8 = [Mg2+] [fear ::8 - 0-210M 3:40M 1-9 = 0.061